# Task-Circuit Quantization: Leveraging Knowledge Localization and Interpretability for Compression

### v1

@@ 1 Introduction
We evaluate TaCQ’s ability to compress models to 2- to 3-bit size on multiple-choice question-answering benchmarks such as MMLU (Hendrycks et al., 2021), a math benchmark, GSM8k (Cobbe et al., 2021), as well as Spider (Yu et al., 2019), a standard text-to-SQL generation benchmark. We see improvements across all settings compared to state-of-the-art mixed-precision PTQ baselines like SqueezeLLM (Kim et al., 2024), SPQR (Dettmers et al., 2023), and SliM-LLM (Huang et al., 2024b) using the same conditioning data and at matched memory budgets. TaCQ especially excels in 2-bit settings, where we improve accuracy over the strongest baseline, SliM-LLM (Huang et al., 2024b), by 15.99% (absolute) on GSM8k, 21.92% on Spider, and 14.35% on MMLU. Moreover, TaCQ also provides consistent improvements in 3-bit settings, and we show that its improvements are consistent across different quantization budgets. Furthermore, while TaCQ obtains the largest gains over baselines in the task-specific setting (i.e. where all methods are conditioned on task-specific data), it also outperforms baselines when using general-purpose pretraining data for calibration, obtaining a $7.20\%$ gain on the MMLU Humanities Split over the strongest baseline SliM-LLM in 2-bits. We also find that previous mixed-precision methods require searching for thresholds to specify bit-width (Dettmers et al., 2023; Kim et al., 2024) or face difficulty adjusting bit-width in a fine grained manner (Huang et al., 2024b). Our method only has one hyperparameter, the fraction of outliers to preserve in 16-bits, which allows us to specify the desired compression ratio deterministically.
@@ 2 Background and Related Work
where $F$ denotes the set of unquantized weights. We focus on this type of training-free post-training quantization for its efficiency and widespread adoption, further providing a complete derivation and descriptions of other quantization techniques in Appendix C.
@@
where $\theta_{\text{normal}}=\theta\setminus\theta_{\text{outliers}}$. Previous approaches have defined weight importance based on local layer-wise loss and weight magnitude (Dettmers et al., 2023; Huang et al., 2024b), other methods ignore the effect of quantization (Shao et al., 2024b; Kim et al., 2024), both fail to predict changes in the global loss due to quantization. In Section 3, we introduce our approach to computing weight sensitivity and in Section 4 we compare our proposed approach with SqueezeLLM, and SPQR, two SOTA approaches that utilize outlier preservation, as well as SliM-LLM (Huang et al., 2024b), a strong mixed-precision baseline which dynamically assigns bit-width based on importance, finding our method significantly outperforms all baselines. Appendix C contains an expanded description of other mixed-precision methods and our choice of baselines.
@@ 3.3 Details on Quantization Implementation
TaCQ uses GPTQ (Frantar et al., 2023a) for quantization, exempting salient weights identified through our metric from consideration. To obtain the gradient term of TaCQ we perform backpropagation from the standard cross-entropy loss. We include in Appendix D a longer specification for our configuration of GPTQ, engineering changes needed to enable downstream conditioning, and a memory-efficient method for computing gradients. We also show that TaCQ is economical in terms of quantization time in Appendix A and that we can reduce the number of examples used to compute gradients while minimally impacting performance in Section G.4.
@@ 4 Results and Analysis
We base our main experimental results and analysis on the Llama-3-8B-Instruct model, and include additional results on Qwen2.5-7B-Instruct, perplexity results on the non-instruction-tuned Llama-3-8B base model, and scaling experiments on Qwen2.5-32B-Instruct in Appendix G.
@@
We include Spider (Yu et al., 2019) as a separate generation-based text-to-SQL task in a zero-shot setting, to simulate a practical application of LLMs. We draw calibration samples from the training set and evaluate on the development set using test-suite accuracy (Zhong et al., 2020). Further details are provided in Appendix E.
@@ 4.1 Main Results
We report our main results for Llama-3-8B-Instruct in Tables 1 and 2. TaCQ shows major improvements over existing SOTA approaches and the GPTQ baseline. In the 2-bit setting, TaCQ surpasses SliM-LLM by absolute margins of 16.0% (20.1% to 36.1%) on GSM8k, 14.1% (34.8% to 49.2%) on MMLU, and 21.9% (0% to 21.9%) on Spider, while other baselines such as GPTQ, SqueezeLLM, and SPQR degrade to near-random performance. In the 3-bit scenario, TaCQ preserves approximately 91%, 96%, and 89% of the unquantized accuracy on GSM8k, MMLU, and Spider, respectively. Moreover, it outperforms the strongest baseline, SliM-LLM, by 1–2% across most datasets.
@@
In Fig. 3, we demonstrate the trade-off between accuracy and average bit-width for different approaches, showing TaCQ consistently outperforms. We describe the hyperparameters tuned for this plot in Appendix A. Additionally, we include a comparison with SPQR in an apples-to-apples setting due to its larger hyperparameter space in Appendix F.
For generation settings where the model must output multiple tokens sequentially, our method represents a qualitative change in the ability of the quantized model. In the Spider task, TaCQ is the only method that recovers non-negligible performance (above zero) in the 2-bit setting. We include qualitative examples in Appendix B for both Spider and GSM8k, where other methods produce random tokens or lose their instruction following capability.
@@ 4.2 Task-Conditioning Transfer
Existing methods assume that quantization performance is relatively unaffected by the choice of calibration set (Williams & Aletras, 2024; Paglieri et al., 2024), and we show in Section D.5 that most methods are implemented such that conditioning on downstream datasets is difficult. We present evidence that calibration datasets have a large impact in 2- to 3-bit settings in Table 3 and Table 4. When conditioning on WikiText2, a standard calibration dataset, SliM-LLM fails to produce any result on MMLU Humanities and reduces to random performance in 2-bit, but conditioning consistently raises performance by more than 10 percentage points. Analyzing the impact for GPTQ at a broader range of bit-widths shows that the effect of conditioning is much smaller at 4 bits and above, and that testing on generation settings can widen the gap.
@@ Appendix A Configurations and Hyperparameter Tuning
We utilize the default configuration in the official implementation: groupsize 16, quantization statistics in 3-bits, and double quantization groupsize of 16. SPQR has 6 adjustable hyperparameters, which gives SPQR a great flexibility in terms of being able to make different tradeoffs. We consider a standard setting in the main tables and include a pareto plot on an apples to apples setting in Appendix F. Specifically for the main tables, for the 3 bit setting we use base bit-width 2, groupsize 16, quantization statistics in 3-bits and double quantization groupsize of 16. For the 2 bit setting we use base bit-width 1, groupsize 16, quantization statistics in 3-bits and double quantization groupsize of 16. The small groupsize of SPQR means that to match bit-width with other methods, it is often compared such that it uses $N-1$ as the base bit-width, where $N$ is the target bit-width (Dettmers et al., 2023; Kim et al., 2024). The number of outliers is variable due to our use of calibration dataset seeds, conditioning datasets, and SPQR’s usage of a threshold. Specifically, the percentage of outliers depends on the calibration dataset seed, model, and bit-width. Critically, as it depends on the conditioning dataset seed, bit-width varies with every run in our main tables. In practice, we observe SPQR to use around 1-2% outliers and we only consider quantization runs where SPQR uses higher average bit-width than our method, noting the lowest bit-width in the bit-width sections of tables. We note that in the non-task conditioned setting with one seed, the thresholding problem is not as big of an issue, as we only need to quantize the model once. Variability arises since we need to condition on each seed and downstream dataset.
For the Pareto plot in Appendix F, we tune the threshold value to increase and decrease the percentage of outliers preserved.
@@
Appendix B Qualitative Results
@@
In Table 7 we show GSM8k CoT generations from all methods in a 2-bit setting for Llama-3-8B-Instruct. We note that while SPQR Dettmers et al. (2023) is the only method not to generate text resembling a response, this is potentially due to the fact that we evaluate SPQR with base-bitwidth of 1 in order to match the approximately 2-bit setting. With higher base bit-width and a different choice of hyperparameters it should be possible for SPQR to achieve better results, for this reason we include the apples to apples setting in Appendix F to marginalize the impact of hyperparameters. In Table 8, we show Spider generation results from all methods in a 2-bit setting for Llama-3-8B-Instruct. TaCQ is the only method to recover task relevant results in 2-bits for Llama-3-8B-Instruct. The next-best method, SliM-LLM, losses instruction following ability, but it does produce legible text. The examples in Table 8 are generally representative of the type of output for each method. For TaCQ we observe that all results look like SQL queries, but not all queries are syntactically or semantically correct. For other methods, the majority of generations do not resemble SQL queries.
Appendix C Expanded Background
@@
Appendix D Expanded Description of TaCQ
@@
D.1 Quantization-Aware Localization (QAL)
@@
D.2 Magnitude-Sharpened Gradient (MSG)
@@
Utilizing the same strategy to replace activations with weights as in Section D.1 we formulate the magnitude-sharpened gradient (MSG) for a specific weight $W_{ij}$ in a single linear layer as
@@
D.3 Computing the Gradient Efficiently.
@@
This approach ensures that the GPU is never responsible for more than the memory cost of a single forward pass, plus a small overhead for the cached activations, if gradient check-pointing is not used to recompute activations. For the experiments reported in 7-8B scale models, all gradient-related computations were performed in 32-bit, though 16-bit gradients are equally possible. We employ the 16-bit strategy for our experiments with the Qwen2.5-32B-Instruct model in Section G.3.
@@
D.4 More details on quantization implementation.
@@
D.5 Inducing Task Conditioning
@@
Appendix E Further Details on Evaluation and Conditioning Datasets
@@
Appendix F SPQR Pareto Comparison
@@
Appendix G Additional Results
G.1 Qwen2.5-7B-Instruct Results
@@ Appendix G Additional Results
G.2 Llama-3-8B Perplexity Results
2-bit – Qwen2.5-32B-Instruct
Method Avg bits MMLU Full
Unquantized 16.00 82.46
SliM-LLM 2.125 62.51
TaCQ 2.10 72.51
@@ G.2 Llama-3-8B Perplexity Results
G.3 Model Size Generalization Test
@@
G.4 Number of Datapoints for Gradient Capture

### v2

@@ 1 Introduction
We evaluate TaCQ’s ability to compress models to 2- to 3-bit size on multiple-choice question-answering benchmarks such as MMLU (Hendrycks et al., 2021), a math benchmark, GSM8k (Cobbe et al., 2021), as well as Spider (Yu et al., 2019), a standard text-to-SQL generation benchmark. We see improvements across all settings compared to state-of-the-art mixed-precision PTQ baselines like SqueezeLLM (Kim et al., 2024), SPQR (Dettmers et al., 2023), and SliM-LLM (Huang et al., 2024b) using the same conditioning data and at matched memory budgets. TaCQ especially excels in 2-bit settings, where we improve accuracy over the strongest baseline, SliM-LLM (Huang et al., 2024b), by 15.99% (absolute) on GSM8k, 21.92% on Spider, and 14.35% on MMLU. Moreover, TaCQ also provides consistent improvements in 3-bit settings, and we show that its improvements are consistent across different quantization budgets and model sizes. Scaling to 32B models, TaCQ recovers 87.93% of the 16-bit unquantized model’s MMLU performance at 2-bit quantization. Furthermore, while TaCQ obtains the largest gains over baselines in the task-specific setting (i.e. where all methods are conditioned on task-specific data), it also outperforms baselines when using general-purpose pretraining data for calibration, obtaining a $7.20\%$ gain on the MMLU Humanities Split over the strongest baseline SliM-LLM in 2-bits. We also find that previous mixed-precision methods require searching for thresholds to specify bit-width (Dettmers et al., 2023; Kim et al., 2024) or face difficulty adjusting bit-width in a fine grained manner (Huang et al., 2024b). Our method only has one hyperparameter, the fraction of outliers to preserve in 16-bits, which allows us to specify the desired compression ratio deterministically.
@@ 2 Background and Related Work
where $F$ denotes the set of unquantized weights. We focus on this type of training-free post-training quantization for its efficiency and widespread adoption, further providing a complete derivation and descriptions of other quantization techniques in Appendix D.
@@
where $\theta_{\text{normal}}=\theta\setminus\theta_{\text{outliers}}$. Previous approaches have defined weight importance based on local layer-wise loss and weight magnitude (Dettmers et al., 2023; Huang et al., 2024b), other methods ignore the effect of quantization (Shao et al., 2024b; Kim et al., 2024), both fail to predict changes in the global loss due to quantization. In Section 3, we introduce our approach to computing weight sensitivity and in Section 4 we compare our proposed approach with SqueezeLLM, and SPQR, two SOTA approaches that utilize outlier preservation, as well as SliM-LLM (Huang et al., 2024b), a strong mixed-precision baseline which dynamically assigns bit-width based on importance, finding our method significantly outperforms all baselines. Appendix D contains an expanded description of other mixed-precision methods and our choice of baselines.
@@ 3.3 Details on Quantization Implementation
TaCQ uses GPTQ (Frantar et al., 2023a) for quantization, exempting salient weights identified through our metric from consideration. To obtain the gradient term of TaCQ we perform backpropagation from the standard cross-entropy loss. We include in Appendix E a longer specification for our configuration of GPTQ, engineering changes needed to enable downstream conditioning, and a memory-efficient method for computing gradients. We also show that TaCQ is economical in terms of quantization time in Appendix A and that we can reduce the number of examples used to compute gradients while minimally impacting performance in Section H.4.
@@ 4 Results and Analysis
We base our main experimental results and analysis on the Llama-3-8B-Instruct model, and include additional results on Qwen2.5-7B-Instruct, perplexity results on the non-instruction-tuned Llama-3-8B base model, and scaling experiments on Qwen2.5-32B-Instruct in Appendix H.
@@
We include Spider (Yu et al., 2019) as a separate generation-based text-to-SQL task in a zero-shot setting, to simulate a practical application of LLMs. We draw calibration samples from the training set and evaluate on the development set using test-suite accuracy (Zhong et al., 2020). Further details are provided in Appendix F.
@@ 4.1 Main Results
2-bit – Qwen2.5-32B-Instruct
Method Avg bits MMLU Full
Unquantized 16.00 82.46
GPTQ 2.00 50.55
SliM-LLM 2.125 62.51
TaCQ 2.10 72.51
Table 1: Application of TaCQ on Qwen2.5-32B-Instruct shows that TaCQ scales to larger models.
@@
We report our main results for Llama-3-8B-Instruct in Tables 2 and 3. TaCQ shows major improvements over existing SOTA approaches and the GPTQ baseline. In the 2-bit setting, TaCQ surpasses SliM-LLM by absolute margins of 16.0% (20.1% to 36.1%) on GSM8k, 14.1% (34.8% to 49.2%) on MMLU, and 21.9% (0% to 21.9%) on Spider, while other baselines degrade to near-random performance. In the 3-bit scenario, TaCQ preserves approximately 91%, 96%, and 89% of the unquantized accuracy on GSM8k, MMLU, and Spider, respectively. Moreover, it outperforms the strongest baseline, SliM-LLM, by 1–2% across most datasets. Additionally, we show that TaCQ generalizes to larger models by quantizing the Qwen2.5-32B-Instruct model, presenting our results in Table 1. TaCQ recovers 87.93% of the 16-bit unquantized model’s performance at 2-bit, reducing the model size by around 8x and outperforming SliM-LLM by 10.00%.
@@
In Fig. 3, we demonstrate the trade-off between accuracy and average bit-width for different approaches, showing TaCQ consistently outperforms. We describe the hyperparameters tuned for this plot in Appendix A. Additionally, we include a comparison with SPQR in an apples-to-apples setting due to its larger hyperparameter space in Appendix G.
For generation settings where the model must output multiple tokens sequentially, our method represents a qualitative change in the ability of the quantized model. In the Spider task, TaCQ is the only method that recovers non-negligible performance (above zero) in the 2-bit setting. We include qualitative examples in Appendix C for both Spider and GSM8k, where other methods produce random tokens or lose their instruction following capability.
@@ 4.2 Task-Conditioning Transfer
Existing methods assume that quantization performance is relatively unaffected by the choice of calibration set (Williams & Aletras, 2024; Paglieri et al., 2024), and we show in Section E.5 that most methods are implemented such that conditioning on downstream datasets is difficult. We present evidence that calibration datasets have a large impact in 2- to 3-bit settings in Table 4 and Table 5. When conditioning on WikiText2, a standard calibration dataset, SliM-LLM fails to produce any result on MMLU Humanities and reduces to random performance in 2-bit, but conditioning consistently raises performance by more than 10 percentage points. Analyzing the impact for GPTQ at a broader range of bit-widths shows that the effect of conditioning is much smaller at 4 bits and above, and that testing on generation settings can widen the gap.
@@ Appendix A Configurations and Hyperparameter Tuning
We utilize the default configuration in the official implementation: groupsize 16, quantization statistics in 3-bits, and double quantization groupsize of 16. SPQR has 6 adjustable hyperparameters, which gives SPQR a great flexibility in terms of being able to make different tradeoffs. We consider a standard setting in the main tables and include a pareto plot on an apples to apples setting in Appendix G. Specifically for the main tables, for the 3 bit setting we use base bit-width 2, groupsize 16, quantization statistics in 3-bits and double quantization groupsize of 16. For the 2 bit setting we use base bit-width 1, groupsize 16, quantization statistics in 3-bits and double quantization groupsize of 16. The small groupsize of SPQR means that to match bit-width with other methods, it is often compared such that it uses $N-1$ as the base bit-width, where $N$ is the target bit-width (Dettmers et al., 2023; Kim et al., 2024). The number of outliers is variable due to our use of calibration dataset seeds, conditioning datasets, and SPQR’s usage of a threshold. Specifically, the percentage of outliers depends on the calibration dataset seed, model, and bit-width. Critically, as it depends on the conditioning dataset seed, bit-width varies with every run in our main tables. In practice, we observe SPQR to use around 1-2% outliers and we only consider quantization runs where SPQR uses higher average bit-width than our method, noting the lowest bit-width in the bit-width sections of tables. We note that in the non-task conditioned setting with one seed, the thresholding problem is not as big of an issue, as we only need to quantize the model once. Variability arises since we need to condition on each seed and downstream dataset.
For the Pareto plot in Appendix G, we tune the threshold value to increase and decrease the percentage of outliers preserved.
@@
Appendix B Quantization Time and Compute Cost
While gradient capture is expensive, there are several reasons why we believe it is tractable in the context of TACQ.
  • Gradients only need to be captured once to freely quantize at varied % outliers and base bitness.
  • Compared to the quantization timelines and overheads of other methods, TACQ has reasonable and sometimes superior quantization timeline, see Table 8.
  • The number of datapoints used for gradient capture can be reduced by 8x without affecting performance. See Table 16.
Method Compute Time (hours) GPU Memory Peak (GB)
GPTQ 7B 0.58 7.68
SLiMLLM 7B 14.66 12.48
TaCQ 7B 2.76 $+$ 0.58 93.29
GPTQ 32B 2.56 20.16
SLiMLLM 32B 54.71 22.608
TaCQ 32B 4.37 $+$ 2.56 138.62
Table 8: Compute time and peak GPU memory for different quantization methods
Appendix C Qualitative Results
@@
In Table 9 we show GSM8k CoT generations from all methods in a 2-bit setting for Llama-3-8B-Instruct. We note that while SPQR Dettmers et al. (2023) is the only method not to generate text resembling a response, this is potentially due to the fact that we evaluate SPQR with base-bitwidth of 1 in order to match the approximately 2-bit setting. With higher base bit-width and a different choice of hyperparameters it should be possible for SPQR to achieve better results, for this reason we include the apples to apples setting in Appendix G to marginalize the impact of hyperparameters. In Table 10, we show Spider generation results from all methods in a 2-bit setting for Llama-3-8B-Instruct. TaCQ is the only method to recover task relevant results in 2-bits for Llama-3-8B-Instruct. The next-best method, SliM-LLM, losses instruction following ability, but it does produce legible text. The examples in Table 10 are generally representative of the type of output for each method. For TaCQ we observe that all results look like SQL queries, but not all queries are syntactically or semantically correct. For other methods, the majority of generations do not resemble SQL queries.
Appendix D Expanded Background
@@
Appendix E Expanded Description of TaCQ
@@
E.1 Quantization-Aware Localization (QAL)
@@
E.2 Magnitude-Sharpened Gradient (MSG)
@@
Utilizing the same strategy to replace activations with weights as in Section E.1 we formulate the magnitude-sharpened gradient (MSG) for a specific weight $W_{ij}$ in a single linear layer as
@@
Concretely, we expect that when QAL is used, the average distance to gridlines will increase. Intuitively, given the same gradient importance, it is more valuable to preserve outliers far away from gridlines as they will be more heavily impacted when quantized. We compute this average distance metric and show that this is indeed the case.
Scale Gradient QAL
Outlier distance to gridlines 3.426e-3 ± 1.604e-3 9.577e-3 ± 4.643e-3
Table 11: Average distance to gridlines for preserved weights under QAL vs gradient selection strategies.
E.3 Computing the Gradient Efficiently.
@@
This approach ensures that the GPU is never responsible for more than the memory cost of a single forward pass, plus a small overhead for the cached activations, if gradient check-pointing is not used to recompute activations. For the experiments reported in 7-8B scale models, all gradient-related computations were performed in 32-bit, though 16-bit gradients are equally possible. We employ the 16-bit strategy for our experiments with the Qwen2.5-32B-Instruct model in Section H.3.
@@
E.4 More details on quantization implementation.
@@
E.5 Inducing Task Conditioning
@@
Appendix F Further Details on Evaluation and Conditioning Datasets
@@
Appendix G SPQR Pareto Comparison
@@
Appendix H Additional Results
H.1 Qwen2.5-7B-Instruct Results
@@ Appendix H Additional Results
H.2 Llama-3-8B Perplexity Results
@@
H.3 Model Size Generalization Test
2-bit – Qwen2.5-32B-Instruct
Method Avg bits MMLU Full
Unquantized 16.00 82.46
GPTQ 2.00 50.55
SliM-LLM 2.125 62.51
TaCQ 2.10 72.51
@@ H.3 Model Size Generalization Test
H.4 Number of Datapoints for Gradient Capture